OpenStudy (anonymous):

helpp mee pleaseeA store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (accidentalaichan):

I remember this Flvs question.....

OpenStudy (anonymous):

lol, exactly what it is i dont know how to do it .

OpenStudy (accidentalaichan):

What are the answer choices?

OpenStudy (anonymous):

there are none

OpenStudy (accidentalaichan):

C = 0.765x + 0.06(0.765x) C=.765*100+.0435*100 C=76.50+4.35 C=$80.85 is the discounted price. ( we replaced x with 100 because x is the current price)

OpenStudy (accidentalaichan):

:) Good luck!

OpenStudy (anonymous):

omg thank you :))

OpenStudy (accidentalaichan):

No problem :)